OpenStudy (amtran_bus):

If A cannot equal B and 1/x+1/A+1/B, then x=?

OpenStudy (anonymous):

What does the equation equal?

OpenStudy (amtran_bus):

Sorry, it is\[A \neq B, \frac{ 1 }{ x }+\frac{ 1 }{ A }=\frac{ 1 }{ }\] Then x=?

OpenStudy (amtran_bus):

that is supposed to be over B

OpenStudy (loser66):



OpenStudy (loser66):

I guide you only. you must do yours. step by step

OpenStudy (loser66):

@Luigi0210

OpenStudy (loser66):

why? i saw you write something, why did you stop?

OpenStudy (amtran_bus):

My internet!

OpenStudy (loser66):

@Luigi0210 take it over, pleaase

OpenStudy (amtran_bus):